OpenStudy (anonymous):

use the quotient rule to differentiate f(x) = (2x + 5) /(2x).

OpenStudy (anonymous):

\[((2)(2x)-(2x+5)(2))/(2x^2)\]

OpenStudy (anonymous):

mistake there: ...../(2x)^2

OpenStudy (anonymous):

can that be simplified any more?

OpenStudy (anonymous):

((2)(2x)−(2x+5)(2))/(2x)^2 =(4x-4x-10)/(4x^2) =-10/(4x^2) =5/2x^2

OpenStudy (anonymous):

5/(2x^2), the x^2 is on the bottom

OpenStudy (anonymous):

ok thank you

OpenStudy (phi):

Just a note: you dropped a minus sign going to the final answer: \[ -\frac{5}{2x^2} \]